# Visual serial processing deficits explain divergences in human and VLM reasoning

### v1

@@ 1 Introduction
To test our hypothesis, we systematically manipulate the serial processing load in three distinct domains. First, in a geometric program induction task inspired by work modeling geometric concepts as programs (Hsu et al., 2022), we vary load by manipulating concept complexity, operationalized as the number of geometric primitives (Minimum Description Length; MDL) required to define a geometric concept (Sablé-Meyer et al., 2022). Second, for visual enumeration, drawing on work distinguishing subitizing from serial counting (Trick & Pylyshyn, 1994), serial load is varied by manipulating color distinctiveness and spatial overlap, known to increase demands on serial attention (Franconeri et al., 2013). Finally, in a mental rotation task based on (Shepard & Metzler, 1971), serial processing demands are modulated by varying the angular disparity between two stimuli, a factor shown to linearly increase human processing time (Cooper & Shepard, 1973). Across these diverse tasks we demonstrate that as task conditions necessitate greater serial processing the VLM-human performance gap widens reliably, providing convergent evidence for a serial processing deficit.
@@ 4 Test Domain 1: Geometric Concepts
The first domain in which we investigate a potential serial processing deficit in VLMs is geometric reasoning. A long tradition in cognitive science has examined how humans reason over geometric patterns, demonstrating that response times (RTs) increase systematically with the structural complexity of the stimulus (Just & Carpenter, 1985). In particular, a growing body of work has argued that humans interpret geometric figures by inferring the underlying generative rules that produced them, rather than merely matching surface features. This program induction approach has been explored in experiments that manipulate the minimum description length (MDL) of symbolic programs used to generate visual stimuli – where MDL reflects the number of discrete compositional steps required to construct the pattern from a fixed vocabulary of geometric primitives and relations (Sablé-Meyer et al., 2022). Such studies suggest that human geometric reasoning is fundamentally compositional and serial: more complex patterns require more sequential operations, resulting in longer RTs but typically preserved accuracy. Recent work introduced the Geoclidean domain-specific language (DSL) to generate 2D geometric stimuli by composing geometric primitives (Hsu et al., 2022). This work revealed that while humans are adept at inferring abstract geometric rules, standard transformer and CNN-based vision encoders are insensitive to this structure. However, it remains unclear whether modern VLMs, which incorporate language-based reasoning, exhibit similar deficits — particularly for complex concepts that demand greater serial processing in humans. We test this by parametrically varying the complexity of geometric concepts generated using the Geoclidean DSL, predicting that model performance will degrade with increasing complexity, even as human accuracy remains stable.
@@ 4.1 Methods
To evaluate the sensitivity of VLMs to geometric concepts, we adapted the “oddball” detection paradigm from foundational studies of non-verbal human geometric intuition (Dehaene et al., 2006). Each trial consisted of an array of six images. Five of these images consistently instantiated a particular geometric concept, while the sixth item—the ’oddball’—violated this concept. We evaluated model and human performance on a geometric oddball task generated using the Geoclidean DSL (Hsu et al., 2022). Stimuli from 37 distinct concepts belonging to two categories — Geoclidean-Elements and Geoclidean-Constraints — were generated. (Figure 1a). These concepts spanned a range of geometric properties, including Euclidean geometry (e.g., parallelism, right angles, line bisecting), specific figures (e.g., squares, triangles, polygons), and constraints and relations (tangency, midpoint relations, angle constraints). We created 100 trials for each of the 37 concepts, yielding a total of 3,700 trials. Human participants each completed a randomly selected subset of 50 trials, drawn from a uniformly sampled 20% subset of the model evaluation trials. These trials were sampled such that each task condition was equally represented, and each trial received at least 20 independent human judgments. The oddball was constructed by removing two relational constraints from the reference concept while preserving its underlying components. Concepts ranged in program complexity from MDL 1 to 4, where Minimum Description Length (MDL) denotes the number of geometric primitives needed to generate the concept. Human participants and VLMs were tasked with identifying the position of the oddball within the array. (Appendix Figure 10). Human participants indicated their choice by clicking on the target image, and their reaction times were recorded. (Appendix Figure 9). VLMs were presented with the same array of images, each overlaid with a red index in the corner, and prompted to output the index corresponding to the oddball. (Appendix A.1.1). We evaluated four large multimodal models — GPT-4o (Achiam et al., 2023), Claude Sonnet 3.7 (Anthropic, 2023), Llama 4 Maverick (AI, 2024), Qwen2.5 VL (Bai et al., 2025) — on this task.
@@ 5 Test Domain 2: Numerical Estimation
Numerical estimation — such as assessing how many objects appear in a scene — is a classic serial processing task. Human numerical estimation abilities are characterized by a transition between the near-instantaneous recognition of small quantities (up to 4–6 items), a process known as subitization, and the serial enumeration of larger quantities (Trick & Pylyshyn, 1994; Kaufman et al., 1949). As humans enumerate larger sets or when visual discrimination becomes challenging, we deploy serial attention to identify each object — binding features to locations and maintaining distinct object representations (Mazza & Caramazza, 2015; Xu & Chun, 2009). Such serial attention-based enumeration becomes more challenging as objects become harder to distinguish (Franconeri et al., 2013; Lavie, 2005). Prior work has documented VLMs’ difficulty in counting (Pothiraj et al., 2025; Campbell et al., 2024b); we predict that, in alignment with human behavior, VLM performance will be highest when the number of objects is low and the objects are non-overlapping. However, while humans maintain accuracy in challenging conditions (Vetter et al., 2008) (e.g. overlapping objects at higher numerosities) by deploying serial attention, we hypothesize that VLM performance will degrade rapidly. Furthermore, we predict that this degradation in performance would coincide with an increase in human reaction time. Critically, we anticipate an asymmetry in performance, where VLM performance increases when overlapping objects are uniquely colored compared to when they are uniformly colored, a pattern not mirrored in human performance. This prediction stems from our hypothesis that VLMs’ serialization capabilities depend heavily on their ability to ground reasoning in their linguistic chain of thought, which is enabled when the objects can be uniquely identified.
@@ 6 Test Domain 3: Mental Rotation
Mental rotation provides a final test case where the sequential nature of the task cannot be effectively substituted with linguistic reasoning. Unlike symbolic pattern recognition, mental rotation requires the manipulation of spatial representations through continuous transformations—operations that are difficult to express in discrete, language-based tokens. A standard mental rotation task presents participants with two images which are either two distinct objects or one object shown from different angles. Decades of cognitive science research demonstrate that humans solve these tasks by mentally rotating one object to align with another, an analog process captured by the robust linear relationship between angular disparity and reaction time (Shepard & Metzler, 1971; Cooper & Shepard, 1973). Functional imaging confirms that this process involves incremental spatial adjustments, engaging parietal regions known to support visuospatial transformation (Zacks et al., 1999; Just & Carpenter, 1985). These transformations rely on what (Kosslyn, 1994) termed “depictive representations” – mental images that preserve geometric structure and spatial relations. Because such analog operations are difficult to serialize into linguistic prompts, mental rotation offers a particularly stringent test of whether VLMs can support visual reasoning that is not scaffolded by language.
@@ 6.1 Methods
We evaluated human and model performance on a classic mental rotation task, adapted from prior cognitive science paradigms (Shepard & Metzler, 1971; Cooper & Shepard, 1973). Stimuli consisted of character-like shapes that were either identical or mirror-reversed versions of a reference, displayed at various rotation angles (Figure 3a). Each trial presented a pair of shapes, and the task was to determine whether they matched or were mirrored. (Appendix Figure 14).
@@ 6.2 Results
We first examined the effect of rotation angle on performance. As shown in Figure 3b, human reaction time increased monotonically with angular disparity, replicating the classic linear RT-angle relationship. This trend reflects the analog nature of human mental rotation: more extreme angles require longer sequences of transformations (Just & Carpenter, 1985; Zacks et al., 1999).
@@ 8 Discussion
In geometric reasoning, as program complexity (measured by MDL) increased, human reaction times rose, whereas VLM performance deteriorated. Previous work using similar tasks argued that human performance implies symbolic program induction, a capability that neural network models lack (Sablé-Meyer et al., 2022). Contrary to this finding, we show that neural networks (here, VLMs) do correlate with both human behavior and measures of task complexity. It is an open question whether to interpret this as evidence that VLMs perform program induction, or that the human behavioral measures do not reflect program induction. However, our results indicate that VLMs appear unable to flexibly compose visual operations in ways that scale with program complexity.
The numerical estimation task revealed that VLMs demonstrate subitizing-like behavior for small quantities under optimal viewing conditions but show performance degradation when object individuation becomes challenging. Humans maintain accuracy in these conditions by deploying serial attention—binding features to locations and maintaining distinct object representations—at the cost of increased processing time (Mazza & Caramazza, 2015; Xu & Chun, 2009). We also find that VLMs benefit from unique coloring in overlapping object conditions, while humans show no comparable advantage. This contrasts with some findings, such as those by (Rahmanzadehgervi et al., 2025), who reported that coloring conditions in numerosity tasks made no significant difference for most of the VLMs tested. Here, we show that when objects overlap, unique coloring provides a distinct and significant advantage to VLMs. This advantage likely arises because distinct colors allow VLMs to individually reference and effectively count objects, thereby enabling them to linguistically ground features and perform feature binding in a way that is not possible in uniformly colored conditions.
Mental rotation provided a particularly stringent test of visual serial processing, as this analog transformation is difficult to decompose into discrete linguistic steps. While humans show a linear relationship between rotation angle and reaction time (Shepard & Metzler, 1971) (reflecting an incremental transformation process), and relatively limited effects of rotation angle on accuracy, VLMs exhibited sharp performance drops as rotation angles increased. This suggests a limitation in performing the continuous transformations that humans execute through depictive representations (Kosslyn, 1994). Importantly, we are not claiming that models are incapable of serial processing—indeed, autoregressive generation implicitly implements serial processing. However, this seems to be a strictly linguistic form of serialization that is difficult to apply to visual tasks like mental rotation.
@@ 8.1 Architectural Improvements
A more promising path forward involves inducing a more sequential and deliberate mode of processing directly, with a new class of models trained using Visually Grounded Reinforcement Learning (RL) showing particular promise. Rather than processing an image in a single pass, methods like ViGoRL, GRIT, UniVG-R1, and Ground-R1 use RL to train a policy that generates an explicit reasoning trace where each textual thought is anchored to a specific coordinate or bounding box in the visual input (Sarch et al., 2025; Wang et al., 2025; Li et al., 2025; Zhao et al., 2025). This process of generating a sequence of region-grounded thoughts functions as a computational analogue to the human saccade-and-fixate cycle, forcing the model to abandon its default holistic strategy and adopt a more structured, pseudo-serial approach to scene analysis. This also increases model interpretability, given the visual-attentional path to an answer can be explicitly examined.
@@ 8.2 Limitations & Future Directions
This work has several limitations. First, our analysis is constrained to a limited set of VLMs. We chose models that are competitive with other available models, but our selection does not necessarily capture the full diversity of performance. Second, our tasks were selected to manipulate serial processing load in visual reasoning; future work could consider a broader range of tasks. Specifically, extending our analysis to tasks like visual search and spatial reasoning (Gilden et al., 2010; Reimer & Schubert, 2019) would increase the generality of our results. Moreover, evaluating spatial reasoning alongside a broader range of tasks with varying serial processing load may allow us to more effectively dissociate spatial reasoning deficits from serial reasoning deficits.

### v2

@@ 1 Introduction
To test our hypothesis, we systematically manipulate the serial processing load in three distinct domains. First, in a geometric program induction task inspired by work modeling geometric concepts as programs (Hsu et al., 2022), we vary load by manipulating concept complexity, operationalized as the number of geometric primitives (Minimum Description Length; MDL) required to define a geometric concept (Sablé-Meyer et al., 2022). Second, for visual enumeration, drawing on work distinguishing subitizing from serial counting (Trick and Pylyshyn, 1994), serial load is varied by manipulating color distinctiveness and spatial overlap, known to increase demands on serial attention (Franconeri et al., 2013). Finally, in a mental rotation task based on (Shepard and Metzler, 1971), serial processing demands are modulated by varying the angular disparity between two stimuli, a factor shown to linearly increase human processing time (Cooper and Shepard, 1973). Across these diverse tasks we demonstrate that as task conditions necessitate greater serial processing the VLM-human performance gap widens reliably, providing convergent evidence for a serial processing deficit.
@@ 4 Test Domain 1: Geometric Concepts
The first domain in which we investigate a potential serial processing deficit in VLMs is geometric reasoning. A long tradition in cognitive science has examined how humans reason over geometric patterns, demonstrating that response times (RTs) increase systematically with the structural complexity of the stimulus (Just and Carpenter, 1985). In particular, a growing body of work has argued that humans interpret geometric figures by inferring the underlying generative rules that produced them, rather than merely matching surface features. This program induction approach has been explored in experiments that manipulate the minimum description length (MDL) of symbolic programs used to generate visual stimuli – where MDL reflects the number of discrete compositional steps required to construct the pattern from a fixed vocabulary of geometric primitives and relations (Sablé-Meyer et al., 2022). Such studies suggest that human geometric reasoning is fundamentally compositional and serial: more complex patterns require more sequential operations, resulting in longer RTs but typically preserved accuracy. Recent work introduced the Geoclidean domain-specific language (DSL) to generate 2D geometric stimuli by composing geometric primitives Hsu et al. (2022). This work revealed that while humans are adept at inferring abstract geometric rules, standard transformer and CNN-based vision encoders are insensitive to this structure. However, it remains unclear whether modern VLMs, which incorporate language-based reasoning, exhibit similar deficits — particularly for complex concepts that demand greater serial processing in humans. We test this by parametrically varying the complexity of geometric concepts generated using the Geoclidean DSL, predicting that model performance will degrade with increasing complexity, even as human accuracy remains stable.
@@ 4.1 Methods
To evaluate the sensitivity of VLMs to geometric concepts, we adapted the “oddball” detection paradigm from foundational studies of non-verbal human geometric intuition (Dehaene et al., 2006). Each trial consisted of an array of six images. Five of these images consistently instantiated a particular geometric concept, while the sixth item—the ’oddball’—violated this concept. We evaluated model and human performance on a geometric oddball task generated using the Geoclidean DSL Hsu et al. (2022). Stimuli from 37 distinct concepts belonging to two categories — Geoclidean-Elements and Geoclidean-Constraints — were generated. (Figure 1a). These concepts spanned a range of geometric properties, including Euclidean geometry (e.g., parallelism, right angles, line bisecting), specific figures (e.g., squares, triangles, polygons), and constraints and relations (tangency, midpoint relations, angle constraints). We created 100 trials for each of the 37 concepts, yielding a total of 3,700 trials. Human participants each completed a randomly selected subset of 50 trials, drawn from a uniformly sampled 20% subset of the model evaluation trials. These trials were sampled such that each task condition was equally represented, and each trial received at least 20 independent human judgments. The oddball was constructed by removing two relational constraints from the reference concept while preserving its underlying components. Concepts ranged in program complexity from MDL 1 to 4, where Minimum Description Length (MDL) denotes the number of geometric primitives needed to generate the concept. Human participants and VLMs were tasked with identifying the position of the oddball within the array. (Appendix Figure 10). Human participants indicated their choice by clicking on the target image, and their reaction times were recorded. (Appendix Figure 9). VLMs were presented with the same array of images, each overlaid with a red index in the corner, and prompted to output the index corresponding to the oddball. (Appendix A.1.1). We evaluated four large multimodal models — GPT-4o (Achiam et al., 2023), Claude Sonnet 3.7 (Anthropic, 2023), Llama 4 Maverick (AI, 2024), Qwen2.5 VL (Bai et al., 2025) — on this task.
@@ 5 Test Domain 2: Numerical Estimation
Numerical estimation — such as assessing how many objects appear in a scene — is a classic serial processing task. Human numerical estimation abilities are characterized by a transition between the near-instantaneous recognition of small quantities (up to 4–6 items), a process known as subitization, and the serial enumeration of larger quantities (Trick and Pylyshyn, 1994; Kaufman et al., 1949). As humans enumerate larger sets or when visual discrimination becomes challenging, we deploy serial attention to identify each object — binding features to locations and maintaining distinct object representations (Mazza and Caramazza, 2015; Xu and Chun, 2009). Such serial attention-based enumeration becomes more challenging as objects become harder to distinguish (Franconeri et al., 2013; Lavie, 2005). Prior work has documented VLMs’ difficulty in counting (Pothiraj et al., 2025; Campbell et al., 2024b); we predict that, in alignment with human behavior, VLM performance will be highest when the number of objects is low and the objects are non-overlapping. However, while humans maintain accuracy in challenging conditions (Vetter et al., 2008) (e.g. overlapping objects at higher numerosities) by deploying serial attention, we hypothesize that VLM performance will degrade rapidly. Furthermore, we predict that this degradation in performance would coincide with an increase in human reaction time. Critically, we anticipate an asymmetry in performance, where VLM performance increases when overlapping objects are uniquely colored compared to when they are uniformly colored, a pattern not mirrored in human performance. This prediction stems from our hypothesis that VLMs’ serialization capabilities depend heavily on their ability to ground reasoning in their linguistic chain of thought, which is enabled when the objects can be uniquely identified.
@@ 6 Test Domain 3: Mental Rotation
Mental rotation provides a final test case where the sequential nature of the task cannot be effectively substituted with linguistic reasoning. Unlike symbolic pattern recognition, mental rotation requires the manipulation of spatial representations through continuous transformations—operations that are difficult to express in discrete, language-based tokens. A standard mental rotation task presents participants with two images which are either two distinct objects or one object shown from different angles. Decades of cognitive science research demonstrate that humans solve these tasks by mentally rotating one object to align with another, an analog process captured by the robust linear relationship between angular disparity and reaction time Shepard and Metzler (1971); Cooper and Shepard (1973). Functional imaging confirms that this process involves incremental spatial adjustments, engaging parietal regions known to support visuospatial transformation Zacks et al. (1999); Just and Carpenter (1985). These transformations rely on what Kosslyn (1994) termed “depictive representations” – mental images that preserve geometric structure and spatial relations. Because such analog operations are difficult to serialize into linguistic prompts, mental rotation offers a particularly stringent test of whether VLMs can support visual reasoning that is not scaffolded by language.
@@ 6.1 Methods
We evaluated human and model performance on a classic mental rotation task, adapted from prior cognitive science paradigms (Shepard and Metzler, 1971; Cooper and Shepard, 1973). Stimuli consisted of character-like shapes that were either identical or mirror-reversed versions of a reference, displayed at various rotation angles (Figure 3a). Each trial presented a pair of shapes, and the task was to determine whether they matched or were mirrored. (Appendix Figure 14).
@@ 6.2 Results
We first examined the effect of rotation angle on performance. As shown in Figure 3b, human reaction time increased monotonically with angular disparity, replicating the classic linear RT-angle relationship. This trend reflects the analog nature of human mental rotation: more extreme angles require longer sequences of transformations (Just and Carpenter, 1985; Zacks et al., 1999).
@@ 8 Discussion
In geometric reasoning, as program complexity (measured by MDL) increased, human reaction times rose, whereas VLM performance deteriorated. Previous work using similar tasks argued that human performance implies symbolic program induction, a capability that neural network models lack Sablé-Meyer et al. (2022). Contrary to this finding, we show that neural networks (here, VLMs) do correlate with both human behavior and measures of task complexity. It is an open question whether to interpret this as evidence that VLMs perform program induction, or that the human behavioral measures do not reflect program induction. However, our results indicate that VLMs appear unable to flexibly compose visual operations in ways that scale with program complexity.
The numerical estimation task revealed that VLMs demonstrate subitizing-like behavior for small quantities under optimal viewing conditions but show performance degradation when object individuation becomes challenging. Humans maintain accuracy in these conditions by deploying serial attention—binding features to locations and maintaining distinct object representations—at the cost of increased processing time Mazza and Caramazza (2015); Xu and Chun (2009). We also find that VLMs benefit from unique coloring in overlapping object conditions, while humans show no comparable advantage. This contrasts with some findings, such as those by Rahmanzadehgervi et al. (2025), who reported that coloring conditions in numerosity tasks made no significant difference for most of the VLMs tested. Here, we show that when objects overlap, unique coloring provides a distinct and significant advantage to VLMs. This advantage likely arises because distinct colors allow VLMs to individually reference and effectively count objects, thereby enabling them to linguistically ground features and perform feature binding in a way that is not possible in uniformly colored conditions.
Mental rotation provided a particularly stringent test of visual serial processing, as this analog transformation is difficult to decompose into discrete linguistic steps. While humans show a linear relationship between rotation angle and reaction time Shepard and Metzler (1971) (reflecting an incremental transformation process), and relatively limited effects of rotation angle on accuracy, VLMs exhibited sharp performance drops as rotation angles increased. This suggests a limitation in performing the continuous transformations that humans execute through depictive representations Kosslyn (1994). Importantly, we are not claiming that models are incapable of serial processing—indeed, autoregressive generation implicitly implements serial processing. However, this seems to be a strictly linguistic form of serialization that is difficult to apply to visual tasks like mental rotation.
@@ 8.1 Architectural Improvements
A more promising path forward involves inducing a more sequential and deliberate mode of processing directly, with a new class of models trained using Visually Grounded Reinforcement Learning (RL) showing particular promise. Rather than processing an image in a single pass, methods like ViGoRL, GRIT, UniVG-R1, and Ground-R1 use RL to train a policy that generates an explicit reasoning trace where each textual thought is anchored to a specific coordinate or bounding box in the visual input (Sarch et al., 2025; Wang et al., 2025; Li et al., 2025; Cao et al., 2025). This process of generating a sequence of region-grounded thoughts functions as a computational analogue to the human saccade-and-fixate cycle, forcing the model to abandon its default holistic strategy and adopt a more structured, pseudo-serial approach to scene analysis. This also increases model interpretability, given the visual-attentional path to an answer can be explicitly examined.
@@ 8.2 Limitations & Future Directions
This work has several limitations. First, our analysis is constrained to a limited set of VLMs. We chose models that are competitive with other available models, but our selection does not necessarily capture the full diversity of performance. Second, our tasks were selected to manipulate serial processing load in visual reasoning; future work could consider a broader range of tasks. Specifically, extending our analysis to tasks like visual search and spatial reasoning Gilden et al. (2010); Reimer and Schubert (2019) would increase the generality of our results. Moreover, evaluating spatial reasoning alongside a broader range of tasks with varying serial processing load may allow us to more effectively dissociate spatial reasoning deficits from serial reasoning deficits.